# Comparison of Nested Geometry Treatments within GPU-Based Monte Carlo Neutron Transport Simulations of Fission Reactors

### v1

@@ 4 Params params)
In ORANGE, this behavior was achieved via a template metaprogramming approach which allows a single set of switch statement logic to be used by all tracking operations, as shown in Listings 2, 3, and 4. In Listing 2, a Traits struct is templated on UType, and template specializations are used to define the corresponding tracker type for each universe type. A visit_universe_type function returns the output of a given functor, which takes a Traits object as an argument. Listing 3 provides a visit_tracker function which returns the output of a supplied functor f, which takes a tracker of arbitrary type as an argument. This is achieved by wrapping f in a second functor—f2, which calls f for a given Traits object—and then calling visit_universe_type with this second functor.
Listing 4 shows how the visit_tracker function can be used to achieve polymorphism. A find_cell functor is first created. This functor, along with a uid and params, is passed visit_tracker, and the resulting cell is returned. Upon compilation, code in Listings 2, 3, and 4 produce bitcode equivalent to that of Listing 1. The other Table 1 tracking operations are implemented in an identical fashion to the find_cell example in Listing 4. As was the case with the DP method, CSG universes in the SP method also use BIH acceleration, which is described in Section 4.3.
@@ 4.3 Single tracker (ST) method
Instead of employing polymorphism, the final approach uses a single CSG tracker for CSG, rectilinear array, and hexagonal array universes. This is accomplished by converting rectilinear and hexagonal array universes into CSG universes. To do so, array cells are explicitly modeled as CSG cells using the method shown in Figure 2. This conversion is done automatically in ORANGE, and the resulting CSG universes are referred to as “pseudo-array universes,” a term specific to this work. Modeling arrays in this fashion is not a new approach: it is the simplest way of modeling an array within an MC code without specialized array universe types.
Because pseudo-array universes are CSG universes, they use the CSG tracking algorithms shown in Section 3.1 rather than the array versions that have improved time complexity. In other words, for pseudo-arrays, find_cell involves an $O(N)$ linear search over all cells and cross_surface requires an $O(N)$ linear search over all neighbor cells. The cross_surface algorithm has an additional complication for pseudo-array universes. In ORANGE, coincident surfaces are always deduplicated to simplify tracking, resulting in a large number of neighbor cells for pseudo-array surfaces. An example of this is shown in Figure 4. In this example, the magenta surface bounds all 32 colored cells. When crossing the magenta surface from the purple cell, the neighbor list includes all 31 gold cells.
@@
To improve the performance of tracking operations for CSG universes, including pseudo-array universes, a BIH acceleration structure was implemented. The BIH construction process is demonstrated in Figure 5. Bounding boxes for each cell are first ascertained. A partition plane is then chosen, and bounding boxes are partitioned into two sets according to the location of each bounding box center. For each of the two sets of partitioned bounding boxes, bounding planes are created by moving the partition to fully enclose all bounding boxes. By performing this process recursively on each set of bounding boxes, a binary tree structure is created, with edges specifying the half-spaces which contain all children.
@@
One key feature of BIH trees is that for a given node, the half spaces created by the bounding planes may overlap. This guarantees that each cell appears in the BIH tree exactly once, unlike $k$-D trees that must store a cell twice if its bounding box is bisected by a partition. This provides an advantage in the pseudo-array use case in which adjacent cells may have bounding boxes that overlap slightly because of floating point error. BIH trees can handle this case without significant performance consequences. BVH trees can also handle this overlap case. However, BVH trees may require more memory because they store six planes per node, significantly more than the two planes required by BIH trees. An open research question within the BIH construction process is how to choose partitions. For the purposes of this work, a standard surface area heuristic (SAH) partitioning scheme (Wald, 2007) was used. This was implemented by evaluating three equally spaced candidate partitions per axes at each partitioning step. Candidate partitions were evaluated using a cost function balancing the number of bounding boxes and the total surfaces area of bounding boxes appearing on each side of the partition.
The find_cell and cross_surface functions are accelerated by traversing the BIH tree. For find_cell, traversal is terminated when a cell is found that contains the supplied point. For cross_surface, the traversal is terminated when a cell is found that contains the supplied point, excluding the cell in which the neutron originated. When traversing the BIH tree, both edge conditions must be tested at each node because the half spaces are allowed to overlap. BIH trees are enabled for all CSG universes—not just pseudo-array universes. As mentioned in Sections 4.1 and 4.2, BIH trees are enabled for all CSG universes in the DP and SP methods as well.
@@ 5 Hardware description
Frontier has 9408 compute nodes, each consisting of a 64-core AMD $\text{3}^{\text{rd}}$ Gen EPYC CPU, and four AMD Radeon Instinct MI250X, each consisting of a two GCDs. Eight CPU cores are reserved for system tasks, leaving 56 usable CPU cores per node. On the CPU side, Frontier also has 512 GB of RAM, but each MI250X has 128 GB of RAM. As a result, Frontier has four times as much GPU RAM per GCD compared to Summit (64 GB vs. 16 GB). The results presented here were compiled with ROCm 5.6.0. It is noted that large variations in performance (up to 50%) where observed as a function of compiler version.
@@ 6 Performance testing
The performance of the three multi-universe GPU tracking methods described in Section 4 relative to RTK (described in Section 3.3) was assessed by obtaining timing results for a rectilinear-array-based reactor problem. A full-core model of the NuScale SMR (Smith, 2017), shown in Figure 6, was chosen for this purpose. The small size of the the NuScale design—about one-eighth the size of a typical PWR—is economically attractive because of its low capital cost (Black et al., 2019), and also permits detailed full-core MC analysis. Likewise, this problem served as the challenge problem for the ExaSMR project within the Exascale Computing Project, in which Shift was used for GPU-based coupled neutron transport / thermal hydraulics analysis (Merzari et al., 2023).
The NuScale design consists of 37 assemblies arranged in a rectilinear grid. Each assembly is a 17$\times$17 array of pins and contains uranium dioxide fuel with a 235U enrichment of either 1.6%, 2.4%, or 3.1% (by mass). The inner circle of 3.1% enriched fuel assemblies contain borosilicate glass burnable neutron absorber rods. Spacer grids and nozzles have been homogenized into slabs for simplicity. For this analysis, the fresh (i.e., non-depleted) fuel version of the problem was used. Fresh fuel contains many fewer nuclides than depleted fuel, thereby minimizing time required to calculate cross sections and maximizing the relative time spent on tracking operations. As a result, inactive cycles with this problem (where no time is spent on tallies) represent the scenario in which tracking operations are expected to comprise the largest fraction of runtime.
@@
For each of the four tracking methods, performance testing was conducted by measuring the tracking rate, i.e., the number of histories simulated per unit time, over a sweep of workloads. Here, workload refers to the number of histories per cycle assigned to each GPU GCD. Each trial consisted of 10 inactive cycles and 10 active cycles using all available GPU GCDs on a single node of Summit or Frontier. During active cycles, the neutron flux was was tallied on a $119\times 119\times 30$ superimposed rectilinear mesh (425,830 mesh volume elements).
@@
Tracking rate results are shown in Figure 7, noting that both plots have different $x$ scales but the same $y$ scale. As expected, the tracking rate increased with workload for all four methods, because large workloads allow the GPU to more effectively hide the latency of memory fetches associated with cross section calculations. Inactive cycle tracking rates are higher because no time is spent processing tally results. On Frontier, tracking rates for all four methods are near their asymptotic limit at a workload of $2.8\times 10^{7}$ histories/cycle/GCD, well before running out of memory, which was found to occur at $5\times 10^{7}$ histories/cycle/GCD for RTK and $3.7\times 10^{7}$ histories/cycle/GCD for the DP, SP, and ST methods. However, on Summit, which has 4×4$\times$4 × less RAM/GCD, all four methods ran out of RAM prior to nearing an asymptotic limit, with RTK running out memory at $1.2\times 10^{7}$ histories/cycle/GCD and the DP, SP, and ST methods running out of memory at $8\times 10^{6}$ histories/cycle/GCD.
Aside from the $1\times 10^{6}$ histories/s/GCD inactive cycles, in which the low workloads cause the kernel launch overhead to wash out any differences between the methods, the relationships between the performance of the four methods remained constant over all workloads. The highest tracking rates were achieved with RTK, as expected, followed by the SP method, the ST method, and the DP method. Further analysis was performed with the $8\times 10^{6}$ histories/cycle/GCD trials on Summit and $2.8\times 10^{7}$ histories/cycle/GCD trials on Frontier. Tracking rates for each of the four methods are shown in Table 2. Results show that all three experimental methods achieve over 90% of the RTK tracking rate on Summit. On Frontier, the DP, SP, and ST methods achieve 77.8%, 91.2%, and 83.4% of the tracking rate during inactive cycles, with all methods achieving at least 87.9% of the RTK tracking rate during active cycles. Therefore, whereas the SP method consistently provides the best performance, no method incurred a significant performance penalty.
@@
(neutron/s/GCD) rate (%) (neutron/s/GCD) rate (%)
@@
The percentages of the total GPU runtime (inactive and active cycles) spent in tracking kernels are shown in Table 3. These values vary from 12.1–26.1%, indicating that tracking operations do not dominate runtime, as expected. Likewise, much larger differences in the performance of the four methods are observed when only considering the time spent within tracking kernels. Figure 8 shows the total time spent within the five principle tracking kernels introduced in Table 2. For a fair comparison, both Summit and Frontier results are from workloads of $8\times 10^{6}$ histories/cycle/GCD. This figure shows that the DP, SP, and ST methods spent 1.70$\times$, 1.14$\times$, and 1.50$\times$ more time in tracking kernels compared to RTK on Summit, and 2.45$\times$, 1.5$\times$, and 2.02$\times$ more time compared to RTK on Frontier. These results highlight the supremacy of the SP method, almost achieving parity with RTK on Summit.
@@
Figure 8: Total time spent in the 5 principle tracking kernels on Summit and Frontier for the NuScale SMR problem, both with $8\times 10^{6}$ histories/cycle/GCD. All tracking kernels not listed here accounted for less than 0.3% of the total tracking time.
Whereas overall tracking rates were slightly higher on Frontier, significantly more time was spent in tracking kernels on Frontier compared to Summit. Frontier also lags behind Summit when comparing GPU performance to CPU performance. Table 4 compares GPU to CPU tracking rates in terms of (1) the number of CPU cores required to match a single GPU GCD, and (2) the speedup, defined as the ratio of the total GPU tracking rate to the total CPU tracking rate on a single node. Noting that both GPU and CPU resources differ between these two machines, Summit has higher CPU core equivalence and speedups compared to Frontier in all cases. This can be attributed to the fact that significantly better CPU performance is achieved on Frontier, where a tracking rate of $4.94\times 10^{3}$ neutron/s/core was achieved for inactive cycles, compared to $2.82\times 10^{3}$ neutrons/s/core for inactive cycles on Summit.
Table 4: GPU to CPU performance comparison for the NuScale SMR problem, run on Summit with $8\times 10^{6}$ histories/cycle/GCD and Frontier with $2.8\times 10^{7}$ histories/cycle/GCD. Speedup refers to the reduction in time-to-solution from using all GPUs on a node compared to using all CPUs.

### v2

@@ 4 Params params)
In ORANGE, this behavior is achieved via a template metaprogramming approach which allows a single set of switch statement logic to be used by all tracking operations, as shown in Listings 2, 3, and 4. In Listing 2, a Traits struct is templated on UType, and template specializations are used to define the corresponding tracker type for each universe type. A visit_universe_type function returns the output of a given functor, which takes a Traits object as an argument. Listing 3 provides a visit_tracker function which returns the output of a supplied functor f, which takes a tracker of arbitrary type as an argument. This is achieved by wrapping f in a second functor—f2, which calls f for a given Traits object—and then calling visit_universe_type with this second functor.
Listing 4 shows how the visit_tracker function can be used to achieve polymorphism. A find_cell functor is first created. This functor, along with a uid and params, is passed to visit_tracker, and the resulting cell is returned. Upon compilation, code in Listings 2, 3, and 4 produce bitcode equivalent to that of Listing 1. The other Table 1 tracking operations are implemented in an identical fashion to the find_cell example in Listing 4. As was the case with the DP method, CSG universes in the SP method also use BIH acceleration, which is described in Section 4.3. Alternatives to this template metaprogramming approach include the curiously recurring template pattern (CRTP) or std::variant with std::visit, which would likely perform similarly.
@@ 4.3 Single tracker (ST) method
While the DP and SP methods benefit from tracking algorithms optimized for specific universe types, it is not clear that these benefits offset the potential performance pitfalls of polymorphism, i.e., virtual function calls in the case of the DP method and divergent execution paths in the case of the DP and SP methods. For comparison, instead of employing multiple polymorphic tracker types, the final approach uses a single CSG tracker for CSG, rectilinear array, and hexagonal array universes. This is accomplished by converting rectilinear and hexagonal array universes into CSG universes. To do so, array cells are explicitly modeled as CSG cells using the method shown in Figure 2. This conversion is done automatically in ORANGE, and the resulting CSG universes are referred to as “pseudo-array universes,” a term specific to this work. Modeling arrays in this fashion is not a new approach: it is the simplest way of modeling an array within an MC code without specialized array universe types.
Because pseudo-array universes are CSG universes, they use CSG tracking algorithms rather than the array versions that have improved time complexity. In other words, for pseudo-arrays, naive implementations of find_cell and cross_surface would each use linear searches over all cells or all neighbors, respectively. The cost of the simple neighbor-based approach is exacerbated by the large number of neighbor cells for pseudo-array surfaces, since planar surfaces in CSG geometries are unbounded. Figure 4 demonstrates the potential inefficiency: when crossing from the purple cell through the magenta surface, which bounds all 32 colored cells, the neighbor list includes all 31 gold cells. It is noted that this issue does not arise with cell-based neighbor lists, an alternative approach in which a mapping between each cell and its neighbor cells is generated dynamically during runtime (Harper et al., 2020).
@@
To improve the performance of tracking operations for CSG universes, including pseudo-array universes, a BIH acceleration structure was implemented. The BIH construction process is demonstrated in Figure 5. First, axes-aligned bounding boxes are generated for each cell using a simple method that involves successively truncating the universe bounding box using the bounding planes associated with each of the cell’s surfaces. Although the nascent implementation of this method in ORANGE does not yet guarantee minimum bounding boxes for arbitrarily complex cells, it is effective for the simple geometric shapes found within the reactor models in this work. Once bounding boxes are ascertained, a partition plane is then chosen, and bounding boxes are partitioned into two sets according to the location of each bounding box center. For each of the two sets of partitioned bounding boxes, bounding planes are created by moving the partition to fully enclose all bounding boxes. By performing this process recursively on each set of bounding boxes, a binary tree structure is created, with edges specifying the half-spaces which contain all children.
@@
One key feature of BIH trees is that for a given node, the half-spaces created by the bounding planes may overlap. This guarantees that each cell appears in the BIH tree exactly once, unlike $k$-D trees that must store a cell twice if its bounding box is bisected by a partition. This provides an advantage in the pseudo-array use case in which adjacent cells may have bounding boxes that overlap slightly because of floating point error. BIH trees can handle this case without significant performance consequences. BVH trees can also handle this overlap case. However, BVH trees may require more memory because they store six planes per node, significantly more than the two planes required by BIH trees. An open research question within the BIH construction process is how to choose partitions. For the purposes of this work, a standard surface area heuristic (SAH) partitioning scheme (Wald, 2007) was used. This was implemented by evaluating three equally spaced candidate partitions per axes at each partitioning step. Candidate partitions were evaluated using a cost function balancing the number of bounding boxes and the total surface area of bounding boxes appearing on each side of the partition.
The find_cell and cross_surface functions are accelerated by traversing the BIH tree. For find_cell, traversal is terminated when a cell is found that contains the supplied point. For cross_surface, the traversal is terminated when a cell is found that contains the supplied point, excluding the cell in which the neutron originated. When traversing the BIH tree, both edge conditions must be tested at each node because the half-spaces are allowed to overlap. BIH trees are enabled for all CSG universes—not just pseudo-array universes. As mentioned in Sections 4.1 and 4.2, BIH trees are enabled for all CSG universes in the DP and SP methods as well.
@@ 5 Hardware description
Frontier has 9408 compute nodes, each consisting of a 64-core AMD $\text{3}^{\text{rd}}$ Gen EPYC CPU, and four AMD Radeon Instinct MI250X, each consisting of two GCDs. Eight CPU cores are reserved for system tasks, leaving 56 usable CPU cores per node. On the CPU side, Frontier also has 512 GB of RAM, but each MI250X has 128 GB of RAM. As a result, Frontier has four times as much GPU RAM per GCD compared to Summit (64 GB vs. 16 GB). Within this work, code was compiled on Frontier with ROCm 5.6.0.
@@ 6 Performance testing
The performance of the three multi-universe GPU tracking methods described in Section 4 relative to RTK (described in Section 3.3) was assessed by obtaining timing results for a rectilinear-array-based reactor problem. A full-core model of the NuScale SMR (Smith, 2017), shown in Figure 6, was chosen for this purpose. The small size of the NuScale design—about one-eighth the size of a typical PWR—is economically attractive because of its low capital cost (Black et al., 2019), and also permits detailed full-core MC analysis. Likewise, this problem served as the challenge problem for the ExaSMR project within the Exascale Computing Project, in which Shift was used for GPU-based coupled neutron transport / thermal hydraulics analysis (Merzari et al., 2023).
The NuScale design consists of 37 assemblies arranged in a rectilinear grid. Each assembly is a $\times17$ array of pins and contains uranium dioxide fuel with a 235U enrichment of either 1.6%, 2.4%, or 3.1% (by mass). The 3.1%-enriched fuel assemblies in the inner circle contain borosilicate glass burnable neutron absorber rods. Spacer grids and nozzles have been homogenized into slabs for simplicity. For this analysis, the fresh (i.e., non-depleted) fuel version of the problem was used. Fresh fuel contains many fewer nuclides than depleted fuel, thereby minimizing the time required to calculate cross sections and maximizing the relative time spent on tracking operations. As a result, inactive cycles within this problem (where no time is spent on tallies) represent the scenario in which tracking operations are expected to comprise the largest fraction of runtime.
@@
The computational representation of this reactor takes on several different forms at runtime when testing the four tracking methods. With the RTK method, the entire core consists of a single RTK universe. With the DP and SP methods, the core consists of an array universe with embedded array universes representing assemblies, each containing CSG universe pins. BIH acceleration is used within these CSG pin universes. This is beneficial because each pin consists of a large number of cells ($\sim 10^{2}$); due to the complexity of the model, pins are split into 40 axial regions. With the ST method, the core, assemblies, and pins are all CSG universes, and each benefits from BIH acceleration.
For each of the four tracking methods, performance testing was conducted by measuring the tracking rate, i.e., the number of histories simulated per unit time, over a sweep of workloads. Here, workload refers to the number of histories per cycle assigned to each GPU GCD. Each trial consisted of 10 inactive cycles and 10 active cycles using all available GPU GCDs on a single node of Summit or Frontier. During active cycles, the neutron flux was tallied on a $119\times 119\times 30$ superimposed rectilinear mesh (425,830 mesh volume elements).
@@
6.1 Tracking rate results
Tracking rate results are shown in Figure 7, noting that both plots have different $x$ scales but the same $y$ scale. As expected, the tracking rate increased with workload for all four methods, because large workloads allow the GPU to more effectively hide the latency of memory fetches associated with cross section calculations. Inactive cycle tracking rates are higher because no time is spent processing tally results. On Frontier, tracking rates for all four methods are near their asymptotic limit at a workload of $2.8\times 10^{7}$ histories/cycle/GCD, well before running out of memory, which was found to occur at $5\times 10^{7}$ histories/cycle/GCD for RTK and $3.7\times 10^{7}$ histories/cycle/GCD for the DP, SP, and ST methods. However, on Summit, which has $4\times$ less RAM/GCD, all four methods ran out of memory prior to nearing an asymptotic limit, with RTK running out of memory at $1.2\times 10^{7}$ histories/cycle/GCD and the DP, SP, and ST methods running out of memory at $8\times 10^{6}$ histories/cycle/GCD.
Aside from the $1\times 10^{6}$ histories/s/GCD inactive cycles, in which the low workloads cause the kernel launch overhead to wash out any differences between the methods, the relationships between the performance of the four methods remained constant over all workloads. The highest tracking rates were achieved with RTK, as expected, followed by the SP method, the ST method, and the DP method. Further analysis was performed with the $8\times 10^{6}$ histories/cycle/GCD trials on Summit and $2.8\times 10^{7}$ histories/cycle/GCD trials on Frontier. Tracking rates for each of the four methods are shown in Table 2. Results show that all three experimental methods achieve over 90% of the RTK tracking rate on Summit. On Frontier, the DP, SP, and ST methods achieve 77.8%, 91.2%, and 83.4% of the RTK tracking rate, respectively, during inactive cycles, with all methods achieving at least 87.9% of the RTK tracking rate during active cycles. These results show that using universe-specific tracker types provides better performance than using a single tracker type for all universe types, provided that this can be done without virtual function calls. However, although the SP method consistently provides the best performance, no method incurred a significant performance penalty.
@@ 6.1 Tracking rate results
(histories/s/GCD) rate (%) (histories/s/GCD) rate (%)
@@
6.2 Performance of individual tracking kernels
[…]
@@ 6.2 Performance of individual tracking kernels
Figure 8: Total time spent in the 5 principal tracking kernels on Summit and Frontier for the NuScale SMR problem, both with $8\times 10^{6}$ histories/cycle/GCD. All tracking kernels not listed here accounted for less than 0.3% of the total tracking time.
6.3 GPU to CPU performance comparison
Whereas overall tracking rates were slightly higher on Frontier, significantly more time was spent in tracking kernels on Frontier compared to Summit. Frontier also lags behind Summit when comparing GPU performance to CPU performance. Table 4 compares GPU to CPU tracking rates in terms of (1) the number of CPU cores required to match a single GPU GCD, and (2) the speedup, defined as the ratio of the GPU tracking rate using all GPUs on a single node to the CPU tracking rate using all CPU cores on a single node. Noting that both GPU and CPU resources differ between these two machines, Summit has higher CPU core equivalence and speedups compared to Frontier in all cases. This can be attributed to the fact that significantly better CPU performance is achieved on Frontier, where a tracking rate of $4.94\times 10^{3}$ histories/s/core was achieved for inactive cycles, compared to $2.82\times 10^{3}$ histories/s/core for inactive cycles on Summit.
Table 4: GPU to CPU performance comparison for the NuScale SMR problem, run on Summit with $8\times 10^{6}$ histories/cycle/GCD and Frontier with $2.8\times 10^{7}$ histories/cycle/GCD. Speedup is defined as the ratio of the GPU tracking rate using all GPUs on a single node to the CPU tracking rate using all CPU cores on a single node.
@@ 6.3 GPU to CPU performance comparison
6.4 Profiling
In order to better understand the performance differences between the four methods, profiling was done using NVIDIA NSight Compute on a V100 (i.e., a Summit GPU) and AMD rocProf on an MI250X (i.e., a Frontier GPU) using a workload of $8\times 10^{6}$ histories/cycle/GCD. Results appear in Table 5. On both the V100 and MI250X, RTK had the highest tracking rates and also the highest theoretical occupancies for the two most expensive kernels (cross_surface and distance_to_surface). The DP method had the lowest tracking rates on the V100 and MI250X and also the lowest theoretical occupancies for cross_surface and distance_to_surface on the V100. On the V100, all kernels except move_within_cell had higher theoretical occupancy with SP compared to DP, which was confirmed to be reflected proportionally in register usage. The fact that the DP and SP methods are identical other than their polymorphism implementation suggests that the virtual functions themselves increase register usage and therefore decrease occupancy. Achieved occupancy was observed to be strongly correlated with theoretical occupancy and did not provide any further insights. On the MI250X, in contrast to the V100, the DP method had the same theoretical occupancy as SP for all kernels except change_direction.
On both the V100 and MI250X, there were no differences in theoretical occupancy between SP and ST because these two methods were run using the exact same kernels within the same executable. This was done because the SP method still needs the ST code for BIH acceleration within CSG universes. While the ST method does not need to be compiled with the SP code in place, stripping out the relatively small and simple SP code did not result in an increase in the ST tracking rate. Consequently, a single executable was maintained for simplicity.
On the V100, only minor differences were observed in the L1 and L2 cache hit rate, and these were not correlated with tracking rate results. This is expected because in all cases, cache performance is likely governed by cross section lookups occurring within the physics kernel launches interspersed with the geometry kernel launches. The differences in branch efficiency between the methods were also not correlated with tracking rate results. Similar results were obtained on the MI250X, noting that (1) only the L2 hit rate can be obtained with rocProf, and (2) instead of branch efficiency, rocProf provides vector arithmetic logic unit (VALU) utilization, which is correlated with branch efficiency.
Finally, on the V100, the number of warp stalls during instruction fetches correlated strongly with tracking rate results.  [^4] Obtained via the smsp__pcsamp_warps_issue_stalled_no_instructions metric. The DP method resulted in about $3\times$ as many warp stalls within the cross_surface and distance_to_surface kernels, suggesting that virtual function calls put additional pressure on the instruction cache.
Table 5: Profiling results on NVIDIA V100 and AMD MI250X GPUs using a workload of $8\times 10^{6}$ histories/cycle/GCD. Highlighted columns show the metrics correlated with performance results.
V100 MI250X Method Kernel Theoretical L1 L2 Branch Warp stalls Theoretical L2 VALU occupancy hit hit efficiency from occupancy hit utilization (%) rate rate (%) instruction (%) rate (%) (%) (%) fetches (%) (#) RTK cross_surface 75 57.8 71.9 89.9 657 100 70.4 86.5 distance_to_surface 50 58.0 66.5 85.9 326 100 53.7 89.6 change_direction 100 45.0 50.4 100 34 100 82.1 88.4 move_within_cell 100 66.7 58.8 0.0 27 100 63.4 86.3 find_cell 100 53.7 69.7 78.8 398 100 58.1 91.7 DP cross_surface 25 59.9 64.4 97.3 4851 50 69.4 87.6 distance_to_surface 12.5 59.6 67.8 89.6 10,031 50 54.5 89.6 change_direction 25 45.0 52.4 100 70 50 51.8 89.6 move_within_cell 100 72.0 69.9 100 23 100 66.2 84.0 find_cell 25 57.1 65.2 92.9 8796 62.5 58.5 91.7 SP cross_surface 37.5 68.3 72.7 97.9 1153 50 69.0 87.6 distance_to_surface 25 72.4 48.9 89.4 3741 50 54.5 89.6 change_direction 50 44.1 50.3 100 36 100 52.3 89.6 move_within_cell 100 72.0 69.9 100 13 100 65.4 84.0 find_cell 50 63.8 75.4 92.0 3710 62.5 58.4 91.7 ST cross_surface 37.5 70.3 74.1 98.7 6110 50 69.2 89.2 distance_to_surface 25 71.7 59.5 87.8 6523 50 54.4 89.6 change_direction 50 41.3 51.6 100 55 100 51.5 89.6 move_within_cell 100 71.95 69.8 100 45 100 65.7 84.0 find_cell 50 66.2 79.2 94.0 9995 62.5 58.2 91.7
6.5 Verification